2. Discouraged workers following statements correctly describe discouraged workers? (If none of the choices apply, leave all of the checkboxes blank.) Check all Which of the that apply. They have given up on looking for a job. They are dissatisfied with their current jobs and are e considering quitting. They are full-time workers who would like to work only part-time. They are counted as part of the labor force by the Bureau of Labor Statistics. The following table contains U.S. employment information from February 2011. Discouraged Unemployed Civillan Population Employed (Persons) (Persons 16 years old and over) (Persons) (Persons) 1,020,000 139,573,000 13,673,000 238,851,000
The following table contains U.S. employment information from February 2011 Civillan Population Employed Unemployed Discouraged (Persons 16 years old and over) (Persons) (Persons) (Persons) 238,851,000 139,573,000 13,673,000 1,020,000 The Bureau of Labor Statistics has several measures of joblessness in addition to the official unemployment rate. One alternative is the U-4 measure of labor underutilization, which is calculated as follows: Unemployed Discouraged U-4-100x
of lebor underutilization, which is calculated as follows: Unemployed+Diacouraged U-4=100 x r Dacuraged Fill in the folowing table by calculating the official unemployment rate and the U-4 measure of labor underutization Official Unemployment Rate U-4 Measure of Labor Underutillization (Percent) (Percent) The official unemployment rate and the U-4 measure of labor underutilization are two different measures of joblessness in the economy. Excluding discouraged workers from the official unemployment rate may cause the official rate to the true extent of underemployment. Grade it Now Save & Continue Contnue without saving
Answer #1

Discouraged workers

HAVE GIVEN UP ON LOOKING FOR A JOB

Official Unemployment rate = Unemployed / (Employed + Unemployed) = 13,673,000 / (139,573,000 + 13,673,000) X 100
= 13,673,000 / 153,346,000 = 8.92%

U-4 Rate = (Unemployed + Discouraged) / (Labor Force + Discouraged) = (13,673,000 + 1,020,000) / (139,573,000 + 13,673,000 + 1,020,000) X 100
= 14,693,000 /  154,266,000 X 100 = 9.52%

It may cause it to UNDERSTATE the true extent of underemployment.